the attendance at a school concert was 578. admission cost $2.00 for adults and $1.50 for children. the receipts for the concert were $985.00. how many adults and how many children attended the concert?
Answer by checkley77(12569)

A+C=578 OR A=578-C
2A+1.5C=985
2(578-C)+1.5C=985
1156-2C+1.5C=985
-.5C=985-1156
-.5C=-171
C=-171/-.5
C=342 CHILDREN'S TICKETS WERE SOLD.
A=578-342=236 ADULT TICKETS WERE SOLD.
PROOF:
2*236+1.5*342=985
472+513=985
985=985